# I Learn Better If You Speak My Language: Understanding the Superior Performance of Fine-Tuning Large Language Models with LLM-Generated Responses

### v2

@@ Abstract
This paper explores an intriguing observation: fine-tuning a large language model (LLM) with responses generated by a LLM often yields better results than using responses generated by humans. We conduct an in-depth investigation to understand why this occurs. Contrary to the common belief that these instances is simply due to the more detailed nature of LLM-generated content, our study identifies another contributing factor: an LLM is inherently more "familiar" with LLM generated responses. This familiarity is evidenced by lower perplexity before fine-tuning. We design a series of experiments to understand the impact of the "familiarity" and our conclusion reveals that this "familiarity" significantly impacts learning performance. Training with LLM-generated responses not only enhances performance but also helps maintain the model’s capabilities in other tasks after fine-tuning on a specific task.
@@
Xuan Ren xuan.ren@adelaide.edu.au                        Biao Wu biaowu165534@gmail.com                        Lingqiao Liuthanks: Corresponding author. lingqiao.liu@adelaide.edu.au
@@ 1 Introduction
Recent research has demonstrated that a large language model (LLM) can generate training data for another LLM Dai et al. (2023); Edwards et al. (2022); Møller et al. (2024); Guo et al. (2023); Ubani et al. (2023); Agrawal et al. (2023); Xu et al. (2023); Kieser et al. (2023); Taori et al. (2023); Peng et al. (2023); Xu et al. (2024). This approach offers a method for transferring knowledge from a larger model to a smaller one, or for creating supplementary training materials, such as rationales Zhang et al. (2024); Kang et al. (2023); Li et al. (2022), or refined step-by-step reasoning Hsieh et al. (2023); Ho et al. (2022); Magister et al. (2023); Fu et al. (2023).
The quality of data generated by large language models (LLMs) is considered to be very high and conducive to effective trainingLee et al. (2023). The prevailing view often attributes the successful training results of LLM-generated data to the fact that LLMs provide more details than human-annotated data, such as ’chain-of-thought-style’ reasoning processHsieh et al. (2023); Ho et al. (2022); Magister et al. (2023); Fu et al. (2023). However, our experiments have found numerous counterexamples. As shown in Figure 1, target labels that include more details and utilize chain of thought styles do not necessarily result in better training outcomes. Through experiments on various tasks, we have observed that when LLMs directly generate target labels, the training results are often superior to those obtained by transforming ground truth into more detailed, step-by-step training. This indicates that whether the answer generation is executed in the LLM’s own style or is based on a rewritten human style is a key factor affecting training outcomes.
Figure 1: This figure illustrates the training outcomes associated with various data generation methods. As demonstrated by the examples shown, the inclusion of additional details does not necessarily translate into enhanced performance.
This observation motivates us to investigate the underlying reason for this phenomenon. We found that whether the LLM is familiar with the target label is a crucial factor influencing model training.
@@
  • Given the same question, an LLM is inherently familiar with responses generated by itself or other LLMs.
  • an LLM can better adapt to and leverage responses in its familiar styles, leading to improved outcomes.
The following section of this paper will examine the hypothesis through a series of experiments. We will measure ’familiarity,’ conduct ablation studies by creating different variants of LLM-generated responses, and evaluate their performance to understand various factors in the process. This will provide insights into our hypothesis.
@@ 2 Empirical Study Protocols
The purpose of this paper is to understand and investigate the aforementioned phenomenon rather than to propose a new method. Our study’s methodology is based on empirical analysis. We begin by introducing our study protocols, including the datasets and evaluation approaches used.
2.1 Evaluation approaches
@@
Without loss of generality, we denote the training data as question and response pairs $(q,r)$. We will use a subset of the LLaMA2 benchmark Touvron et al. (2023) to create the training dataset. We will use a subset of the LLaMA2 benchmark [34] that contains human-labeled explicit reasoning processes to create the training dataset. The question and response pairs in these datasets are human-constructed. Additionally, we will use an LLM to generate new responses, following $r^{\prime}=LLM(q,P)$, where $LLM$ denotes the language model and $P$ denotes the prompt used to create the response. We will explore different $P$ in the following experiment. This creates a synthetic version of dataset $\{(q,r^{\prime})\}$.
@@
Our training approach for the LLaMA2-13b-chatTouvron et al. (2023) and Mistral-7b-Instruct-v2Jiang et al. (2023) models involves specific adaptations to optimize performance. Notably, the warm-up period is set to 10% of the total training steps consistently across our experiments. For the LLaMA-13b-chat model training on the HumanEval dataset, we set the learning rate to 2e-4. This adjustment was necessary as lower rates proved insufficient for the model to effectively learn from this dataset. Given the small size of the training datasets, such as MBPP and HumanEval, we reduced the batch size to 10 to accommodate the limited data. In contrast, for general fine-tuning of the LLaMA2-13b-chat and Mistral-7b-Instruct-v2 models, we employ a default learning rate of 2e-5, a batch size of 32, and a warm-up period comprising 10% of the total steps. Additionally, we use a cosine learning rate schedule and fine-tune only the Q and V matrices of the LORAHu et al. (2021) parameters with a rank of 8. All models are trained and evaluated using half-precision.
@@
We evaluate the fine-tuned model in two different ways. One is called in-domain evaluation, which assesses performance on the tasks of the training set. The other is called cross-task evaluation, which evaluates performance across tasks on other datasets. The purpose of using cross-domain evaluation is to test if the model can benefit from the training process of a similar task and if the model will forget its capability on other tasks after training on a specific task.
3 Experiment setup
3.1 Dataset and Evaluation
@@
We utilized the LLaMA2 benchmark Touvron et al. (2023) to select five tasks that contain human-labeled explicit reasoning. These were evaluated based on accuracy to assess the effectiveness of different data construction methods. The term ’dataset with explicit reasoning’ refers to datasets that include a detailed reasoning process in their annotations. This category includes GSM8KCobbe et al. (2021), MATHHendrycks et al. (2021), HumanEvalChen et al. (2021), MBPPAustin et al. (2021), and ECQA Aggarwal et al. (2021), where the ground truth comprises a complete human-labeled problem-solving process or rationale, not just the final answer. For example, the ground truth in GSM8K involves comprehensive mathematical reasoning, and the coding datasets’ ground truths consist of complete codes, which are logical sequences by nature. We included coding datasets like MBPP and HumanEval in this category, as developing code inherently involves logical reasoning, even though it may not always be presented in the form of a traditional chain of thought. Our model was trained on one dataset and subjected to in-domain testing on the same dataset, as well as cross-domain testing on other datasets. Our selection includes all coding and mathematics datasets within the LLaMA2 benchmark. We chose to use the ECQA dataset because, unlike the CQATalmor et al. (2018) dataset included in the LLaMA2 benchmark, ECQA includes rationales provided by human annotators.
@@
Code: We evaluated coding ability using the HumanEval Chen et al. (2021) and MBPP Austin et al. (2021) datasets. We report the pass@1 accuracy for each dataset. Note that the HumanEval dataset contains only 164 testing data points with no training data. We divided the training task into two parts: in the first part, we trained on the first 82 data points and tested on the last 82. In the second part, we reversed this order. Ultimately, we averaged the test results from both parts. For the MBPP-full dataset, as opposed to the MBPP-sanitized version which includes fewer training data points, we utilized all available training data (374 data points) and testing data (500 data points)
@@
Evaluation Script:
To facilitate automatic evaluation across all models and datasets, we developed a specialized prompt format requiring models to conclude their predictions with ’Final Answer: answer.’ For instance, the output should read ’Final Answer: 100’ for an answer of 100. Our evaluation script extracts the text following the ’Final Answer:’ substring. For the ECQA task, it verifies if the first text after ’Final Answer:’ correctly matches one of the answer options. For mathematical problems including the GSM8K and MATH datasets, the script checks if the number after the substring ’Final Answer:’ aligns with the ground truth number. We utilize the built-in testing script for the HumanEval dataset to ensure result accuracy and employ our custom scripts for all other tasks. By adapting the training data’s question and answer formats to this style, our models maintain consistent response formatting across different tasks. The evaluation scripts for HumanEval Chen et al. (2021) and MBPP Austin et al. (2021) require answers in a specific format. Models that have not been trained on this format struggle to follow it correctly, even when provided with in-context examples. Consequently, we do not conduct cross-domain evaluations on the MBPP and HumanEval datasets.
4 Observations and Hypothesis:
4.1 Observation 1: Superior Performance of GPT-4 Over Human-Annotated Labels
@@
Groundtruth HumanEval, Mistral 0.3625 0.191 0.583 0.323
@@
GPT-4 Answer Directly 0.1245 0.105 0.552 0.159
@@
Table 1: Comparison of human-annotated data versus data generated directly by GPT-4. We employ GPT-4 to generate responses using only the questions from the training datasets. For the ECQA task, we provide GPT-4 with gold labels (excluding rationales) to guide its response generation. In-domain performance is highlighted in grey. Data points significantly lower in accuracy by 15% or more relative to another method are marked in red.
Our study is inspired by a compelling observation: when a sophisticated LLM, such as GPT-4, is used to generate responses to questions, and these questions along with their generated responses are used as a synthetic dataset, fine-tuning smaller LLMs (we use both Llama2-13B-Chat and Mistral-7B-Instruct-v2) on this synthetic dataset often yields better performance than using the original dataset with human-provided answers. The detailed experimental results are presented in Table 1. As demonstrated, the majority of the low-performance data points (14 out of 15) marked in red reflect accuracies when training with human-annotated ground truth. Using synthetic datasets consistently yields higher in-domain performance compared to human-answered datasets. Notably, we observe a significant improvement in math-related tasks, often exceeding a 10% absolute increase in in-domain performance. Additionally, the model trained on the synthetic data achieves higher cross-domain performance. This observation motivates us to investigate the underlying reasons.
4.2 Observation 2: significantly lower perplexity of GPT-4 and other LLMs Over Human-Annotated Labels
Figure 2: Average Perplexity Comparison
We measure the perplexity of responses as assessed by the target (smaller) LLMs. An interesting observation, shown in Figure 2, is that, given the input questions from the responses generated by GPT-4 attain significantly lower perplexity compared to those provided by humans. To explore whether this phenomenon is unique to GPT-4 or common among LLM responses, we employed a zero-shot approach with the Mistral-7B-Instruct-v2 model and the LLaMA-13B-chat model. These models answered the same questions from the training set. As shown in Figure 2, their perplexity was also significantly lower than that of the human-labeled dataset. Lower perplexity indicates that LLMs are more familiar with responses generated by themselves or other LLMs than with human-generated responses.
4.3 Hypotheses
From Observation 2, we notice that given the same question, the average perplexity of responses generated by LLMs is consistently lower than the perplexity of ground truth annotated by humans. This observation leads us to pose our first hypothesis: When generating target labels for the same question, one LLM inherently more familiar with data generated by another LLM. We not only observed consistently lower perplexity from GPT-4 across tasks in Observation 2 but also noted that LLMs trained on GPT-4 generated responses consistently outperform the same LLMs trained using human-annotated data. This phenomenon leads us to propose our second hypothesis: An LLM performs better when trained on data it is familiar with.
5 Investigation of the proposed hypotheses
Groundtruth GSM8K, Mistral 0.434 0.162 0.594 4.54 128.88
GPT-4 Answer Directly 0.597 0.246 0.597 2.35 179.94
Step by step 0.574 0.224 0.591 2.27 282.9
Step by step transformation of GT 0.499 0.190 0.6 2.67 235.892
Groundtruth GSM8K, Llama 0.364 0.141 0.565 3.27 128.88
GPT-4 Answer Directly 0.428 0.128 0.575 2.30 179.941
Step by step 0.402 0.134 0.553 2.19 290.92
Step by step transformation of GT 0.394 0.146 0.573 2.43 242.22
Groundtruth Math algebra, Mistral 0.264 0.206 0.554 5.83 163.36
GPT-4 Answer Directly 0.553 0.302 0.608 2.20 328.76
Step by step 0.533 0.239 0.583 2.249 451.99
Step by step transformation of GT 0.444 0.210 0.577 2.65 370.35
Groundtruth Math algebra, Llama 0.36 0.126 0.575 5.15 163.36
GPT-4 Answer Directly 0.35 0.150 0.561 2.59 328.76
Step by step 0.344 0.145 0.593 2.55 450.08
Step by step transformation of GT 0.376 0.141 0.565 2.81 369.06
Groundtruth ECQA, Mistral 0.258 0.134 0.68 51.3 73.51
GPT-4 Answer Directly 0.462 0.223 0.722 6.35 176.54
Step by step 0.481 0.203 0.71 4.66 325.15
Step by step transformation of GT 0.462 0.194 0.681 6.89 218.748
Groundtruth ECQA, Llama 0.132 0.0798 0.631 11.96 73.51
GPT-4 Answer Directly 0.379 0.156 0.656 5.33 176.54
Step by step 0.363 0.140 0.648 4.46 337.30
Step by step transformation of GT 0.312 0.108 0.628 5.55 226.99

### v3

@@ Abstract
This paper explores an intriguing observation: fine-tuning a large language model (LLM) with responses generated by a LLM often yields better results than using responses generated by humans, particularly in reasoning tasks. We conduct an in-depth investigation to understand why this occurs. Contrary to the common belief that these instances is due to the more detailed nature of LLM-generated content, our study identifies another contributing factor: an LLM is inherently more "familiar" with LLM generated responses. This familiarity is evidenced by lower perplexity before fine-tuning. We design a series of experiments to understand the impact of the "familiarity" and our conclusion reveals that this "familiarity" significantly impacts learning performance. Training with LLM-generated responses not only enhances performance but also helps maintain the model’s capabilities in other reasoning tasks after fine-tuning on a specific task. Our code and data are public at https://github.com/XuanRen4470/I-Learn-Better-If-You-Speak-My-Language.
@@
Xuan Ren University of Adelaide xuan.ren@adelaide.edu.au                        Biao Wu University of Technology Sydney biaowu165534@gmail.com                        Lingqiao Liuthanks:   Corresponding author. University of Adelaide lingqiao.liu@adelaide.edu.au
@@ 1 Introduction
Recent research has demonstrated that a large language model (LLM) can generate training data for another LLM Dai et al. (2023); Edwards et al. (2022); Møller et al. (2024); Guo et al. (2023); Ubani et al. (2023); Piedboeuf and Langlais (2023a); Agrawal et al. (2023); Xu et al. (2023); Kieser et al. (2023); Taori et al. (2023); Peng et al. (2023); Xu et al. (2024). This approach offers a method for transferring knowledge from a larger model to a smaller one, or for creating supplementary training materials, such as rationales Zhang et al. (2024); Kang et al. (2023); Li et al. (2022), or refined step-by-step reasoning Hsieh et al. (2023); Ho et al. (2022); Magister et al. (2023); Fu et al. (2023).
In this paper, we made an intriguing observation that fine-tuning an LLM (dubbed target LLM) with responses generated by a LLM, either target LLM or other LLMs, often yields better results than using responses provided by humans in chain-of-thought reasoning tasks, as shown in Table 1. The prevailing view attributes the successful training results of LLM-generated data to the fact that LLMs provide more details than human-annotated data, such as ‘chain-of-thought-style’ reasoning processHsieh et al. (2023); Ho et al. (2022); Magister et al. (2023); Fu et al. (2023). However, our experiments have found numerous counterexamples. As shown in Figure 1 , target responses that include more details and utilize chain of thought styles do not necessarily result in better training outcomes. This indicates there are other factors contributing to the excellent performance of LLM generated responses.
Figure 1: This figure shows training outcomes for different data generation methods, demonstrating that more details do not always improve performance.
The observation above motivates us to investigate the underlying reason for this phenomenon. We found that whether the target LLM is familiar with the target responses is a crucial factor influencing model training on reasoning tasks.
@@
  • Given the same question, the target LLM is inherently familiar with responses generated by itself or other LLMs.
  • The target LLM can better adapt to and leverage responses in its familiar styles, leading to improved outcomes.
The following section of this paper will examine the hypothesis through a series of experiments. We will measure ‘familiarity,’ conduct ablation studies by creating different variants of LLM-generated responses, and evaluate their performance to understand various factors in the process. This will provide insights into our hypothesis.
@@ 2 Empirical Study Protocols
The purpose of this paper is to understand and investigate the aforementioned phenomenon through empirical analysis. We begin by introducing our study protocols, including the datasets and evaluation approaches used.
2.1 Training and Evaluation approaches
@@
Without loss of generality, we denote the training data as question and response pairs $(q,r)$. We will use a subset of the Llama2 benchmark Touvron et al. (2023) to create the training dataset. We will use a subset of the Llama2 benchmark that contains human-labeled explicit reasoning processes to create the training dataset. The question and response pairs in these datasets are human-constructed. Additionally, we will use an LLM to generate new responses, following $r^{\prime}=LLM(q,P)$, where $LLM$ denotes the language model and $P$ denotes the prompt used to create the response. We will explore different $P$ in the following experiment. This creates a synthetic version of dataset $\{(q,r^{\prime})\}$.
@@
Our training approach for the Llama2-13b-chatTouvron et al. (2023) and Mistral-7b-Instruct-v2Jiang et al. (2023) models involves specific adaptations to optimize performance. For general, fully supervised fine-tuning of the Llama2-13b-chat and Mistral-7b-Instruct-v2 models, we employ a default learning rate of 2e-5, a batch size of 32, and a warm-up period comprising 10% of the total steps. In contrast, Given the small size of the training datasets, such as MBPP and HumanEval, we reduced the batch size to 10 to accommodate the limited data. For the Llama2-13b-chat model training on the HumanEval dataset, we set the learning rate to 2e-4. This adjustment was necessary as lower rates proved insufficient for the model to effectively learn from this dataset. Additionally, we use a cosine learning rate schedule and fine-tune only the Q and V matrices of the LORAHu et al. (2021) parameters with a rank of 8. All models are trained and evaluated using half-precision.
@@
We evaluate the fine-tuned model in two different ways. One is called in-domain evaluation, which assesses performance on the tasks of the training set. The other is called cross-task evaluation, which evaluates performance across tasks on other datasets. The purpose of using cross-task evaluation is to test if the model can benefit from the training process of a similar task and if the model will forget its capability on other tasks after training on a specific task.
We have developed an evaluation script to automatically measure and report the pass@1 accuracy of models for each dataset. For the HumanEval dataset, we utilize its built-in testing script to ensure the accuracy of results. Our evaluation scripts require answers to follow a specific format, detailed in Appendix F. The models not trained in this format have demonstrated difficulty adhering to it for coding tasks, such as MBPP and HumanEval Chen et al. (2021); Austin et al. (2021). Therefore, cross-domain evaluations are not conducted for MBPP and HumanEval.
2.2 Dataset and API
@@
We utilized the Llama2 benchmark Touvron et al. (2023) to select five datasets that include a detailed reasoning process in their annotations. This category includes GSM8KCobbe et al. (2021), MATHHendrycks et al. (2021), HumanEvalChen et al. (2021), MBPPAustin et al. (2021), and ECQA Aggarwal et al. (2021), where the ground truth comprises a complete human-labeled problem-solving process or rationale, not just the final answer. For example, the ground truth in GSM8K involves comprehensive mathematical reasoning, and the coding datasets’ ground truths consist of complete codes, which are logical sequences by nature. We included coding datasets like MBPP and HumanEval in this category, as developing code inherently involves logical reasoning, even though it may not always be presented in the form of a traditional chain of thought. Our model was trained on one dataset and subjected to in-domain testing on the same dataset, as well as cross-domain testing on other datasets. Our selection includes all coding and mathematics datasets within the Llama2 benchmark. We chose to use the ECQA dataset because it includes human annotated rationales, unlike the CQATalmor et al. (2018) dataset included in the Llama2 benchmark.
@@
Code: We evaluated coding ability using the HumanEval Chen et al. (2021) and MBPP Austin et al. (2021) datasets. Note that the HumanEval dataset contains only 164 testing data points with no training data. We divided the training task into two parts: in the first part, we trained on the first 82 data points and tested on the last 82. In the second part, we reversed this order. Ultimately, we averaged the test results from both parts. For the MBPP-full dataset, as opposed to the MBPP-sanitized version which includes fewer training data points, we utilized all available training data (374 data points) and testing data (500 data points)
@@
API: We did our experiment with use gpt-4-1106-preview OpenAI (2023) from OpenAI and claude-3-5-sonnet-20240620 Anthropic (2023) from Anthropic.
3 Observation and Hypothesis
3.1 Observation 1: Superior Performance of LLMs Over Human-Annotated Labels
Our study is inspired by a compelling observation: when an advanced LLM, such as GPT-4 or Claude 3.5, is used to generate responses to questions, and these questions along with their generated responses are used as a synthetic dataset, fine-tuning a smaller LLM (such as Llama2-13B-Chat or Mistral-7B-Instruct-v2) on this synthetic dataset often yields better performance than using the original dataset with human-provided answers. The detailed experimental results are presented in Table 1. As demonstrated, the majority of the low-performance data points (14 out of 17), marked in red, reflect accuracies when training with human-annotated ground truth. Using synthetic datasets consistently yields higher in-domain performance compared to human-answered datasets. Notably, we observe a significant improvement in math-related tasks, often exceeding a 10% absolute increase in in-domain performance. Additionally, the model trained on the synthetic data achieves higher cross-domain performance. This observation motivates us to investigate the underlying reasons.
3.2 Observation 2: Significantly Lower Perplexity of LLM generated responses Over Human-Annotated Labels
Figure 2: Average Perplexity Comparison
We measured the perplexity of responses as assessed by the target LLMs (Mistral-7B-Instruct-v2 and Llama2-13B-chat). An interesting observation, shown in Figure 2, is that responses generated by LLMs such as GPT-4, Claude 3.5, Mistral-7B-Instruct-v2, and Llama2-13B-chat, using a zero-shot approach, exhibit significantly lower perplexity compared to those provided by humans for the same input questions from the training data. Lower perplexity suggests that LLMs are more familiar with responses generated by other LLMs or themselves than with human-generated responses.
3.3 Hypotheses
From Observation 2, we notice that given the same question, the average perplexity of responses generated by LLMs is consistently lower than the perplexity of ground truth annotated by humans. This observation leads us to pose our first hypothesis: When generating target responses for the same question, one LLM inherently more familiar with data generated by another LLM. We not only observed consistently lower perplexity from advanced LLMs (GPT-4 and Claude 3.5) across tasks in Observation 2 but also noted that LLMs trained on advanced LLMs generated responses consistently outperform the same LLMs trained using human-annotated data. This phenomenon leads us to propose our second hypothesis: An LLM performs better when trained on data it is familiar with.
@@ 3.3 Hypotheses
Claude Answer Directly 0.38 0.129 0.678
Groundtruth HumanEval, Mistral 0.363 0.191 0.583 0.323
@@
GPT-4 Answer Directly 0.125 0.105 0.552 0.159
Claude Answer Directly 0.296 0.104 0.435 0.140
@@
Claude Answer Directly 0.339 0.122 0.546 0.204
Table 1: Human-annotated data Vs. data generated directly by GPT-4/Claud 3.5. In-domain performance is highlighted in grey. Data points are highlighted when accuracy is more than 15% below the highest accuracy on the same dataset and model. There are 14, 1, and 2 red data points for Groundtruth, GPT-4, and Claude, respectively.
4 Investigation of the hypotheses
Method Training Dataset and Model Type GSM8K Math Algebra ECQA token length
Zeroshot Mistral 0.413 0.185 0.504
Zeroshot Llama2 0.35 0.125 0.597
GPT-4 Answer Directly Mistral 0.597 0.246 0.597 179.94
GPT-4 Step-by-step 0.574 0.224 0.591 282.9
GPT-4 Detailed Step-by-Step transformation of GT 0.506 0.196 0.602 315.028
GPT-4 Answer Directly Llama 0.428 0.128 0.575 179.941
GPT-4 Step-by-step 0.402 0.134 0.553 290.92
GPT-4 Detailed Step-by-Step transformation of GT 0.396 0.114 0.553 315.028
GPT-4 Answer Directly Mistral 0.553 0.302 0.608 328.76
GPT-4 Step-by-step 0.533 0.239 0.583 451.99
GPT-4 Detailed Step-by-Step transformation of GT 0.468 0.223 0.582 484.368
GPT-4 Answer Directly Llama 0.35 0.150 0.561 328.76
GPT-4 Step-by-step 0.344 0.145 0.593 450.08
GPT-4 Detailed Step-by-Step transformation of GT 0.299 0.110 0.545 484.368
GPT-4 Answer Directly Mistral 0.462 0.223 0.722 176.54
GPT-4 Step-by-step 0.481 0.203 0.71 325.15
GPT-4 Detailed Step-by-Step transformation of GT 0.487 0.186 0.68 322.855
GPT-4 Answer Directly Llama 0.379 0.156 0.656 176.54
GPT-4 Step-by-step 0.363 0.140 0.648 337.30
GPT-4 Detailed Step-by-Step transformation of GT 0.135 0.106 0.66 322.855